OpenStudy (anonymous):

Suggestion for a small change that will go a long way? I answer a lot of questions in the Mathematics section, so I find myself using the equation tool often. However, theres one little problem that frustrates me often with this tool. Whenever you input something into the equation solver, the page automatically scrolls down to the bottom where your answer is being typed. This makes it very difficult for me to copy equations in the question, because I have to repeatedly scroll up to their question several times to input a small piece of their equation, then I get pushed back down and have to

OpenStudy (anonymous):

...do it all over again. Maybe this sounds like too small of a problem to complain about but it frustrates me daily.

OpenStudy (kinggeorge):

If you're trying to copy what someone used in the equation editor, there's a very simple way. Right click the equations you want to copy, click "Show Math As," and then click "TeX Commands." You can try it out on this.\[\lim_{n\to\infty} \left(1+\frac{1}{n}\right)^n\]

OpenStudy (anonymous):

Usually, the asker won't use the equation tool, and will just type something in plain text in their question, so I can't copy it without looking at it and typing it myself into the equation tool

OpenStudy (kinggeorge):

You could temporarily copy/paste what they wrote into your response so you have a reference that's easily visible. Just remember to delete it before you post.

OpenStudy (anonymous):

Yes, I guess thats a decent "solution" but it just seems like something that shouldn't be necessary to me